OpenStudy (anonymous):

A mural measure 115 feet long by 255 feet high. Find the ratio of the height to the perimeter of this mural

OpenStudy (anonymous):

Assuming the mural is rectangular, then its perimeter is 2(lenth) + 2(width) = 2(115) + 2(255) = 230 + 510 = 740 So the ratio of height to perimeter is 255:740

OpenStudy (anonymous):

Explain how you got 2? There is a formula with it?